# Most incompatible measurements and sum-of-squares optimisation

### v1

@@ I Introduction
Important progress has been made in the case of pairs of measurements. In DFK19 , the authors characterised the most incompatible pairs of measurement in a fixed dimension. Through the relation to steering, these results yield a one-sided device-independent dimension witness, which has even been demonstrated experimentally DSU+21 . In a complementary direction, TFR+21 considered the scenario where the number of outcomes is fixed instead of the dimension, and identified the most incompatible pairs in that setting. Together, these works provide a rather complete picture for measurement incompatibility when restricted to pairs, albeit from different but complementary perspectives.
@@ II Preliminaries
We start by fixing some notations used in this work: $k$ will always be the number of measurements, $d$ the dimension of the underlying Hilbert space, and $n_{1},\ldots,n_{k}$ the numbers of outcomes. As for the indices, $x\in\{1,\ldots,k\}$ will label the measurements and $a\in\{1,\ldots,n_{c}\}$ the outcome, so that a set of measurements will be denoted by $\{\{A_{a|x}\}_{a}\}_{x}$, often abbreviated $\{A_{a|x}\}$. Denoting $\delta$ the Kronecker delta and ${\vec{\jmath}}=j_{1},\ldots,j_{k}$, we can state the definition of compatibility, also called joint measurability. A set $\{A_{a|x}\}$ of measurements is jointly measurable if it admits a parent measurement, that is, a measurement $\{G_{\vec{\jmath}}\}_{\vec{\jmath}}$ such that, for all $a$ and $x$,
@@
Note that this is a slight change of notation compared to Des22 : the quantity $H^{\mathrm{g}}_{k,d}$ defined there is denoted $H^{\mathrm{g}}_{k}(d)$ here. Finally, for convenience, we write $\chi^{\ast}_{k}(n)=\chi^{\ast}(n,\ldots,n)$ when all $k$ numbers of outcomes are identical.
III Degree two
@@
III.1 Random incompatibility robustness
@@
Note that the value of $\alpha_{k,n}$ chosen initially is precisely picked to maximise the resulting value of $\eta$. For $k=2$, this result surpasses the best known bound for $\eta^{\mathrm{r}}$, namely, $\frac{1}{2}[1+1/(n+1)]$ (DFK19, , Eq. (35)) (setting $n_{A}=n_{B}=n$). One could in principle use the same technique to make this bound dependent on different numbers of outcomes.
III.2 Application to anticommuting observables
@@
Note that the case $k=2$ was already known, see (TFR+21, , Appendix B.4). To the best of our knowledge, such results are new, although they were to be expected as similar observations have been obtained in the context of entropic uncertainty relations WW08 . They may give rise to a semi-device independent certification of the number of dichotomic measurements as studied in VPLU24 .
On the one hand, $\chi^{\mathrm{r}}_{k}(2)\geqslant\frac{1}{\sqrt{k}}$ directly comes from Eq.˜21 when setting $n=2$, and this can be extended to $\eta^{\mathrm{g}}$ using Eq.˜14. On the other hand, anticommuting observables allow to reach these bounds; such observables being known to exist from dimension $2^{\lfloor\frac{k}{2}\rfloor}$ on Sha00 . This tightness is a direct application of the upper bounds from (DFK19, , Eq. (E6)), which read:
@@
III.3 Depolarising incompatibility robustness
For $\eta^{\mathrm{d}}$ (as well as $\eta^{\mathrm{g}}$), we can exploit the post-processing monotonicity of the incompatibility measure to study only rank-one measurements, which are extremal DFK19 . The results obtained this way are naturally agnostic of the number of outcomes. This section follows closely Section˜III.1, trading the number of outcomes for the dimension; this parallel will be generalised in Section˜V.2 below.
@@ III.3 Depolarising incompatibility robustness
where $\alpha_{k,d}$ is defined in Eq.˜17 and the proportionality factor is again fixed by normalisation. Our goal is to find the noise associated with this $H_{\vec{\jmath}}^{(2)}$, that is, the value of $\eta$ such that we have a feasible point in Eq.˜5. This requires finding the marginals and the normalisation factor of $H_{\vec{\jmath}}^{(2)}\propto t_{\vec{\jmath}}T_{\vec{\jmath}}^{2}-2\alpha_{k,d}t_{\vec{\jmath}}T_{\vec{\jmath}}+\alpha_{k,d}^{2}t_{\vec{\jmath}}\mathds{1}$:
@@
$\sum_{\vec{\jmath}}t_{\vec{\jmath}}T_{\vec{\jmath}}$ $=kd^{k-1}\mathds{1},$ $\sum_{\vec{\jmath}}\delta_{j_{x},a}t_{\vec{\jmath}}T_{\vec{\jmath}}$ $=d^{k-1}P_{a|x}+(k-1)d^{k-2}\mathds{1},$ (26)
$\sum_{\vec{\jmath}}t_{\vec{\jmath}}T_{\vec{\jmath}}^{2}$ $=k(d+k-1)d^{k-2}\mathds{1},$ $\sum_{\vec{\jmath}}\delta_{j_{x},a}t_{\vec{\jmath}}T_{\vec{\jmath}}^{2}$ $=\big{(}d+2(k-1)\big{)}d^{k-2}P_{a|x}+(k-1)(d+k-2)d^{k-3}\mathds{1}.$ (27)
@@
IV Degree three
In this section, we keep the form of the universal parent measurement (as a polynomial in $S_{\vec{\jmath}}$) but climb one step in the degree ladder. For a fixed number of outcomes, we extend Section˜III.1 and obtain tighter bounds on $\eta^{\mathrm{g}}$. For a fixed dimension, we extend Section˜III.3 and give consequences of the obtained bounds on $\eta^{\mathrm{g}}$ for the experimental observation of genuine high-dimensional steering.
IV.1 Fixed number of outcomes
Like in Section˜III.1, we consider $k$ projective measurements with $n$ outcomes $\{P_{j_{1}|1}\}_{j_{1}=1}^{n},\ldots,\{P_{j_{k}|k}\}_{j_{k}=1}^{n}$, extending the final result to all measurements with the pre-processing monotonicity of $\eta^{\mathrm{g}}$. With
@@ IV.1 Fixed number of outcomes
The linear and quadratic terms have already been tackled in Eqs.˜19 and 20, and we can use $\mathds{1}\geqslant P_{a|x}$ to obtain
@@
Note that the value of $\beta_{k,n}$ chosen initially is precisely picked to maximise the resulting value of $\eta$. For $n=2$, we obtain again the value studied in Section˜III.2, which is no surprise as this value, being already optimal, could not be improved on. For $k=2$, this result reproduces (TFR+21, , Proposition B.12) (setting $n_{A}=n_{B}=n$).
IV.2 Fixed dimension
Like in Section˜III.3, we consider $k$ rank-one measurements in dimension $d$, denoted $\{A_{j_{1}|1}\}_{j_{1}},\ldots,\{A_{j_{k}|k}\}_{j_{k}}$, extending the final result to all measurements by using the post-processing monotonicity of $\eta^{\mathrm{g}}$. We write again $A_{j_{x}|x}=(\operatorname{Tr}A_{j_{x}|x})Q_{j_{x}|x}$ with $Q_{j_{x}|x}$ rank-one projectors. With this, we recall that
@@ IV.2 Fixed dimension
where $\beta_{k,d}$ is defined in Eq.˜30 and the proportionality factor is fixed by normalisation. Note that $H_{\vec{\jmath}}^{(3)}$ is semidefinite positive since the polynomial $X(X-\beta_{k,n})^{2}$ takes nonnegative values on the interval $[0,k]$ containing the eigenvalues of $T_{\vec{\jmath}}$. Like in Section˜IV.1, our goal is to find the noise associated with this $H_{\vec{\jmath}}^{(3)}$. This requires finding a lower bound on the marginals of $H_{\vec{\jmath}}^{(3)}\propto t_{\vec{\jmath}}T_{\vec{\jmath}}^{3}-2\beta_{k,d}t_{\vec{\jmath}}S_{\vec{\jmath}}^{2}+\beta_{k,d}^{2}t_{\vec{\jmath}}S_{\vec{\jmath}}$.
The linear and quadratic terms were treated in Eqs.˜26 and 27, and we can use $(\operatorname{Tr}A_{a|x})\mathds{1}\geqslant A_{a|x}$ to obtain
@@
and for its marginals, we need to use, instead of Eq.˜33, the inequality
@@
IV.3 Application to genuine high-dimensional steering
Now that we have dimension-dependent bounds, we can derive semi device-independent dimension witnesses DSU+21 . In this work we do not define genuine high-dimensional steering and refer to Des22 for details. For $d$-preparable assemblages, plugging Eq.˜42 into (Des22, , Eq. (9)), we have that
@@ IV.3 Application to genuine high-dimensional steering
for which we give numerical values in Table˜1a. Note that Eq.˜43 is symmetric by exchange of $k$ and $d$; this only appears to us to be a coincidence.
From Eq.˜43 we can derive the desired witness, namely, an upper bound on the dimension of the system that depends on the (measurable) steering robustness. Denoting $\mathrm{SR}=\mathrm{SR}_{\{\sigma_{a|x}\}}$ for conciseness, we have
@@
$\leavevmode\nobreak\ k\leavevmode\nobreak\$$\leavevmode\nobreak\ d\leavevmode\nobreak\$2 3 4 5 6
@@
$\leavevmode\nobreak\ k\leavevmode\nobreak\$$\leavevmode\nobreak\ d\leavevmode\nobreak\$2 3 4 5 6
@@
Table 1: Upper bound on the steering robustness for $d$-preparable assemblages obtained with $k$ settings. (a) Analytical values obtained with Eq.˜43. The bold values indicate the cases for which we could strictly improve on the state of the art, the other values matching it; see (Des22, , Table I) for comparison. For $k=2$, our result indeed reproduces (Des22, , Eq. (25)), while for $k=d+2$, it coincides with the cloning machine bound (Des22, , Eq. (11)). (b) Numerical values obtained by solving the SDP (52) at level $t=3$. Bold values show cases where we outperform Table˜1a.
V Higher orders: unveiling a hierarchy
@@
V.1 Degree four and mutually unbiased bases
We consider rank-one projective measurements onto $k$ MUBs and use the notations from Section˜III.1, where $n=d$. We refer to DEBZ10 for the definition of MUBs, standard constructions, and applications. Here we only need their defining property: $P_{j_{x_{1}}|x_{1}}P_{j_{x_{2}}|x_{2}}P_{j_{x_{1}}|x_{1}}=\frac{1}{d}P_{j_{x_{1}}|x_{1}}$ for $x_{1}\neq x_{2}$. This property allows us to rewrite $P_{j_{x_{1}}|x_{1}}P_{j_{x_{2}}|x_{2}}P_{j_{x_{1}}|x_{1}}P_{j_{x_{2}}|x_{2}}=\frac{1}{d}P_{j_{x_{1}}|x_{1}}P_{j_{x_{2}}|x_{2}}$, so that $S_{\vec{\jmath}}^{4}$ can be normalised.

### v2

@@ I Introduction
Important progress has been made in the case of pairs of measurements BHSS13 . In DFK19 , the authors characterised the most incompatible pairs of measurement in a fixed dimension. Through the relation to steering, these results yield a one-sided device-independent dimension witness, which has been demonstrated experimentally DSU+21 . In a complementary direction, TFR+21 considered the scenario where the number of outcomes is fixed instead of the dimension, and identified the most incompatible pairs in that setting. Together, these works provide a rather complete picture for measurement incompatibility when restricted to pairs, albeit from different but complementary perspectives.
@@ II Preliminaries
We start by fixing some notations used in this work: $k$ will always be the number of measurements, $d$ the dimension of the underlying Hilbert space, and $n_{1},\ldots,n_{k}$ the numbers of outcomes. As for the indices, $x\in\{1,\ldots,k\}$ will label the measurements and $a\in\{1,\ldots,n_{x}\}$ the outcome, so that a set of measurements will be denoted by $\{\{A_{a|x}\}_{a}\}_{x}$, often abbreviated $\{A_{a|x}\}$. Denoting $\delta$ the Kronecker delta and ${\vec{\jmath}}=j_{1},\ldots,j_{k}$, we can state the definition of compatibility, also called joint measurability. A set $\{A_{a|x}\}$ of measurements is jointly measurable if it admits a parent measurement, that is, a measurement $\{G_{\vec{\jmath}}\}_{\vec{\jmath}}$ such that, for all $a$ and $x$,
@@
Note that this is a slight change of notation compared to Des22 : the quantity $H^{\mathrm{g}}_{k,d}$ defined there is denoted $H^{\mathrm{g}}_{k}(d)$ here. Finally, for convenience, we write $\chi^{\ast}_{k}(d;n)=\chi^{\ast}(d;n,\ldots,n)$ and $\chi^{\ast}_{k}(n)=\chi^{\ast}(n,\ldots,n)$ when all $k$ numbers of outcomes are identical.
III Previous works
In this section we summarise some known results on the topic, sometimes rewriting them in our notation.
We start with BQG+17 , the only published numerical results on the topic that we are aware of. Therein, an extensive study is conducted, with a focus on qubit measurements. Note that the authors were not aware that different incompatibility measures could give rise to different answers to the question of the most incompatible measurements, so that the distinction between $\eta^{\mathrm{d}}$ and $\eta^{\mathrm{g}}$ is not very apparent in the text, although it was used in the code for performance.
In HMZ16 , the symmetric approximate cloning machine from KW99 is used to obtain
$H^{\mathrm{d}}_{k}(d)\geqslant\frac{1}{k}\left(1+\frac{k-1}{d+1}\right),\quad\text{thus}\quad H^{\mathrm{g}}_{k}(d)\geqslant\frac{1}{k}\left(1+2\ \frac{k-1}{d+1}\right)$ (17)
using Eq.˜14. The idea is to approximately clone the input state into $k$ copies which are then fed to the measurements. By duality, the (depolarising) noise can be transferred to the measurements.
In BN18 , the following result is derived in the framework of free spectrahedra:
$\chi^{\mathrm{r}}_{k}(2)=\frac{1}{\sqrt{k}},$ (18)
which is attained by projective measurements coming from $k$ anticommuting observables. Also there, the dimension-dependent bound $\chi^{\mathrm{r}}_{k}(d;2)\geqslant 1/(2d)$ is derived, but we do not highlight it here as it was later improved, see Eq.˜22 below.
In DFK19 , universal parent POVMs are explicitly given to show that
$H^{\mathrm{g}}_{2}(d)\geqslant\frac{d-2+\sqrt{d^{2}+4d-4}}{4(d-1)},\quad\chi^{\mathrm{r}}(n_{1},n_{2})\geqslant\frac{1}{2}\left(1+\frac{1}{\sqrt{n_{1}n_{2}}+1}\right),\quad\text{and}\quad H^{\mathrm{g}}_{2}(d)=\frac{1}{2}\left(1+\frac{1}{\sqrt{d}}\right).$ (19)
In (BN20, , Proposition 6.7), the cloning bound is extended to the random incompatibility robustness:
$\chi^{\mathrm{r}}(d;n_{1},\ldots,n_{k})\geqslant\frac{1}{k}\left(1+\frac{k-1}{d\max_{x}n_{x}+1}\right).$ (20)
In (TFR+21, , Appendix B.4), the technique from DFK19 for fixed dimension is extended to the setting for fixed number of measurements, yielding
$\chi^{\mathrm{g}}_{2}(n_{1},n_{2})\geqslant\frac{1}{2}\left(1+\frac{\sqrt{n_{1}}+\sqrt{n_{2}}+2}{n_{1}+n_{2}+\sqrt{n_{1}}+\sqrt{n_{2}}}\right).$ (21)
In BN22 , the following dimension-dependent bound is proven:
$\chi^{\mathrm{r}}_{k}(d;2)\geqslant 4^{-\lfloor d/2\rfloor}\binom{2\lfloor d/2\rfloor}{\lfloor d/2\rfloor}\sim\sqrt{\frac{2}{\pi d}}$ (22)
asymptotically.
In Des22 , the pairwise optimal bound from DFK19 is used to obtain (loose) bounds for more than two measurements, namely,
$H^{\mathrm{g}}_{k}(d)\geqslant\big(H^{\mathrm{g}}_{2}(d)\big)^{r}\left[1-2\,\Big(1-H^{\mathrm{g}}_{2}(d)\Big)\left(1-\frac{2^{r}}{k}\right)\right],$ (23)
where $r=\lfloor\log_{2}k\rfloor$.
In BEK+25 , methods from non-commutative polynomial optimization are combined with the analysis of the extreme points of some free spectrahedra to demonstrate that
$\chi^{\mathrm{r}}(2,n)\leqslant\frac{1}{2}\left(1+\frac{1}{\sqrt{n}+1}\right)\quad\text{and}\quad\chi^{\mathrm{r}}_{4}(2;2)\leqslant\frac{2}{\sqrt{13}}.$ (24)
The authors then conjecture both inequalities to be tight. Note their notation for $\chi^{\mathrm{r}}(d;n_{1},\ldots,n_{k})$ is $s_{\mathbb{C}}(d,k,(n_{1},\ldots,n_{k}))$.
IV Degree two
@@
IV.1 Random incompatibility robustness
@@
Note that the value of $\alpha_{k,n}$ chosen initially is precisely picked to maximise the resulting value of $\eta$. For $k=2$, this result surpasses the best known bound for $\eta^{\mathrm{r}}$, namely, $\frac{1}{2}[1+1/(n+1)]$, see Eq.˜19 (setting $n_{1}=n_{2}=n$). One could in principle use the same technique to make this bound dependent on different numbers of outcomes.
IV.2 Application to anticommuting observables
@@
These results were already obtained in BN18 , the proof therein being way more geometric than ours given below. They may give rise to a semi-device independent certification of the number of dichotomic measurements as studied in VPLU24 .
On the one hand, $\chi^{\mathrm{r}}_{k}(2)\geqslant\frac{1}{\sqrt{k}}$ directly comes from Eq.˜29 when setting $n=2$, and this can be extended to $\eta^{\mathrm{g}}$ using Eq.˜14. On the other hand, anticommuting observables allow to reach these bounds; such observables being known to exist from dimension $2^{\lfloor\frac{k}{2}\rfloor}$ on New32 . This tightness is a direct application of the upper bounds from (DFK19, , Eq. (E6)), which read:
@@
IV.3 Depolarising incompatibility robustness
For $\eta^{\mathrm{d}}$ (as well as $\eta^{\mathrm{g}}$), we can exploit the post-processing monotonicity of the incompatibility measure to study only rank-one measurements, which are extremal DFK19 . The results obtained this way are naturally agnostic of the number of outcomes. This section follows closely Section˜IV.1, trading the number of outcomes for the dimension; this parallel will be generalised in Section˜VI.2 below.
@@ IV.3 Depolarising incompatibility robustness
where $\alpha_{k,d}$ is defined in Eq.˜25 and the proportionality factor is again fixed by normalisation. Our goal is to find the noise associated with this $H_{\vec{\jmath}}^{(2)}$, that is, the value of $\eta$ such that we have a feasible point in Eq.˜5. This requires finding the marginals and the normalisation factor of $H_{\vec{\jmath}}^{(2)}\propto t_{\vec{\jmath}}T_{\vec{\jmath}}^{2}-2\alpha_{k,d}t_{\vec{\jmath}}T_{\vec{\jmath}}+\alpha_{k,d}^{2}t_{\vec{\jmath}}\mathds{1}$:
@@
$\sum_{\vec{\jmath}}t_{\vec{\jmath}}T_{\vec{\jmath}}$ $=kd^{k-1}\mathds{1},$ $\sum_{\vec{\jmath}}\delta_{j_{x},a}t_{\vec{\jmath}}T_{\vec{\jmath}}$ $=d^{k-1}A_{a|x}+(k-1)d^{k-2}\mathds{1},$ (34)
$\sum_{\vec{\jmath}}t_{\vec{\jmath}}T_{\vec{\jmath}}^{2}$ $=k(d+k-1)d^{k-2}\mathds{1},$ $\sum_{\vec{\jmath}}\delta_{j_{x},a}t_{\vec{\jmath}}T_{\vec{\jmath}}^{2}$ $=\big(d+2(k-1)\big)d^{k-2}A_{a|x}+(k-1)(d+k-2)d^{k-3}\mathds{1}.$ (35)
@@
V Degree three
In this section, we keep the form of the universal parent measurement (as a polynomial in $S_{\vec{\jmath}}$) but climb one step in the degree ladder. For a fixed number of outcomes, we extend Section˜IV.1 and obtain tighter bounds on $\eta^{\mathrm{g}}$. For a fixed dimension, we extend Section˜IV.3 and give consequences of the obtained bounds on $\eta^{\mathrm{g}}$ for the experimental observation of genuine high-dimensional steering.
V.1 Fixed number of outcomes
Like in Section˜IV.1, we consider $k$ projective measurements with $n$ outcomes $\{P_{j_{1}|1}\}_{j_{1}=1}^{n},\ldots,\{P_{j_{k}|k}\}_{j_{k}=1}^{n}$, extending the final result to all measurements with the pre-processing monotonicity of $\eta^{\mathrm{g}}$. With
@@ V.1 Fixed number of outcomes
The linear and quadratic terms have already been tackled in Eqs.˜27 and 28, and we can use $\mathds{1}\geqslant P_{a|x}$ to obtain
@@
Note that the value of $\beta_{k,n}$ chosen initially is precisely picked to maximise the resulting value of $\eta$. For $n=2$, we obtain again the value studied in Section˜IV.2, which is no surprise as this value, being already optimal, could not be improved on. For $k=2$, this result reproduces Eq.˜21 (setting $n_{1}=n_{2}=n$).
V.2 Fixed dimension
Like in Section˜IV.3, we consider $k$ rank-one measurements in dimension $d$, denoted $\{A_{j_{1}|1}\}_{j_{1}},\ldots,\{A_{j_{k}|k}\}_{j_{k}}$, extending the final result to all measurements by using the post-processing monotonicity of $\eta^{\mathrm{g}}$. We write again $A_{j_{x}|x}=(\operatorname{Tr}A_{j_{x}|x})Q_{j_{x}|x}$ with $Q_{j_{x}|x}$ rank-one projectors. With this, we recall that
@@ V.2 Fixed dimension
where $\beta_{k,d}$ is defined in Eq.˜38 and the proportionality factor is fixed by normalisation. Note that $H_{\vec{\jmath}}^{(3)}$ is semidefinite positive since the polynomial $X(X-\beta_{k,n})^{2}$ takes nonnegative values on the interval $[0,k]$ containing the eigenvalues of $T_{\vec{\jmath}}$. Like in Section˜V.1, our goal is to find the noise associated with this $H_{\vec{\jmath}}^{(3)}$. This requires finding a lower bound on the marginals of $H_{\vec{\jmath}}^{(3)}\propto t_{\vec{\jmath}}T_{\vec{\jmath}}^{3}-2\beta_{k,d}t_{\vec{\jmath}}S_{\vec{\jmath}}^{2}+\beta_{k,d}^{2}t_{\vec{\jmath}}S_{\vec{\jmath}}$.
The linear and quadratic terms were treated in Eqs.˜34 and 35, and we can use $(\operatorname{Tr}A_{a|x})\mathds{1}\geqslant A_{a|x}$ to obtain
@@
and for its marginals, we need to use, instead of Eq.˜41, the inequality
@@
V.3 Application to genuine high-dimensional steering
Now that we have dimension-dependent bounds, we can derive semi device-independent dimension witnesses DSU+21 . In this work we do not define genuine high-dimensional steering and refer to Des22 for details. For $d$-preparable assemblages, plugging Eq.˜50 into (Des22, , Eq. (9)), we have that
@@ V.3 Application to genuine high-dimensional steering
for which we give numerical values in Table˜1a. Note that Eq.˜51 is symmetric by exchange of $k$ and $d$; this only appears to us to be a coincidence.
From Eq.˜51 we can derive the desired witness, namely, an upper bound on the dimension of the system that depends on the (measurable) steering robustness. Denoting $\mathrm{SR}=\mathrm{SR}_{\{\sigma_{a|x}\}}$ for conciseness, we have
@@
$\penalty 10000\ k\penalty 10000\$$\penalty 10000\ d\penalty 10000\$2 3 4 5 6
@@
$\penalty 10000\ k\penalty 10000\$$\penalty 10000\ d\penalty 10000\$2 3 4 5 6
@@
Table 1: Upper bound on the steering robustness for $d$-preparable assemblages obtained with $k$ settings. (a) Analytical values obtained with Eq.˜51. The bold values indicate the cases for which we could strictly improve on the state of the art, the other values matching it; see (Des22, , Table I) for comparison. For $k=2$, our result indeed reproduces (Des22, , Eq. (25)), while for $k=d+2$, it coincides with the cloning machine bound (Des22, , Eq. (11)). (b) Numerical values obtained by solving the SDP (60) at level $t=3$. Bold values show cases where we outperform Table˜1a.
VI Higher orders: unveiling a hierarchy
@@
VI.1 Degree four and mutually unbiased bases
We consider rank-one projective measurements onto $k$ MUBs and use the notations from Section˜IV.1, where $n=d$. We refer to DEBZ10 for the definition of MUBs, standard constructions, and applications. Here we only need their defining property: $P_{j_{x_{1}}|x_{1}}P_{j_{x_{2}}|x_{2}}P_{j_{x_{1}}|x_{1}}=\frac{1}{d}P_{j_{x_{1}}|x_{1}}$ for $x_{1}\neq x_{2}$. This property allows us to rewrite $P_{j_{x_{1}}|x_{1}}P_{j_{x_{2}}|x_{2}}P_{j_{x_{1}}|x_{1}}P_{j_{x_{2}}|x_{2}}=\frac{1}{d}P_{j_{x_{1}}|x_{1}}P_{j_{x_{2}}|x_{2}}$, so that $S_{\vec{\jmath}}^{4}$ can be normalised.